A college student earns $5000 by working a full-time job every summer. He invested half the money in a mutual fund and half the money in a stock that yielded four times as much interest as the mutual fund. After a year he earned $250 in interest. What are the interest rates of the mutual fund and the stock?
Answer by addingup(3677)

Let x be the interest from his mutual fund investment. Then, 4x will be the interest from the money in stocks:
2500x+2500(4x)= 250
2500x+10000x = 250
12500x = 250
x = 250/12500 = 0.02 = 2%
And stocks: 4(2) = 8%